# Limiting noise and outliers in the MICROSCOPE space test of the weak equivalence principle

### v3

@@ Abstract
The MICROSCOPE space mission has tested the weak equivalence principle (universality of free fall) finding no violation to $\simeq 10^{-15}$, about two orders of magnitude better than on the ground. We review this result looking for lessons for the future. The acceleration difference of the falling bodies is limited by random thermal noise attributed to losses in the tiny gold wire which connects each of them to the enclosing cage. Our analysis reveals anomalies which are inexplicable within the current best knowledge of such noise. The anomalies may be due to a fluctuating sub-$\mu\rm V$ potential caused by electric charge patches on the surface of the test bodies and the electrodes coupled to large acceleration/voltage biases. Since the acceleration difference is derived from the individual measurements of each test body, it depends upon their arbitrarily defined zero, and this causes the bias. The measurement data show a large number of outliers, acceleration spikes (“glitches”) induced by thermal stress from the Earth and the Sun on the insulating blankets of the spacecraft. They occur at the synodic frequency relative to Earth (also the frequency of a violation signal) and the Sun, their harmonics and their difference. Outliers are removed and the ensuing gaps (up to $35\%$, $40\%$ of the sessions data) are filled with artificial reconstructed data. However, gaps and artificial data retain memory of the outliers they have replaced and may therefore mimic a violation signal or cancel an effect (signal or systematic). An existing alternative approach relying solely on real measured data would make the result of the experiment fully robust and unquestionable. A correct model of the stiffness of the wire, the presence of a fluctuating patch potential noise with the same frequency dependence as thermal noise from internal damping, as well as a thermal insulation that does not give rise to glitches, were all anticipated by the GOCE gravity mission.
@@ I Introduction
The instrument complement included two sensor units (SU), one with test masses (TM) of different composition – Pt and Ti– for the WEP test (SUEP), the other with equal composition Pt TMs for control (SUREF). Each SU includes two test masses configured as coaxial hollow cylinders, the common axis being the sensitive axis for the detection of a potential WEP violation signal. Each TM forms part of an independent accelerometer. A set of electrodes is used for measuring the position of the TM as well as applying the voltages that maintain it “motionless” with respect to its cage. A thin gold wire provides electric grounding and polarization of the electrostatically levitated mass. TM motion is detected by capacitive sensing (displacements induce capacitance variations). In orbit, the axis of symmetry of the SUs is in the orbit plane and a putative violation signal is an Earth pointing vector whose size oscillates at the orbital frequency of $0.16818\,\rm{mHz}$ plus or minus the satellite spin rate. In the valid experiment sessions, two spin rates opposite the orbit motion were used, resulting in the signal frequencies $0.92499\,\rm{mHz}$ and $3.11133\,\rm{mHz}$, the synodic frequencies relative to the Earth in V2 and V3 spin mode respectively.
The measurements were taken in sessions of 120 orbits (8.26 days) or less, the upper limit being driven by periodic correction of clock drift.
Early results published in 2017 MS-PRL2017 were based, for the SUEP sensor, on the analysis of one measurement session (#218) lasting $8.26$ days and reported a fractional difference of $\simeq 10^{-14}$ in the accelerations of Pt and Ti test bodies. A complete analysis of 94 days in 19 sessions reports no violation at the level of $\eta\simeq 10^{-15}$-PRL2022 ; CQGspecialissueMS2022 , an improvement by about two orders of magnitude over the best ground tests with rotating torsion balances Adel2008 ; AdelCQGfocusissue2012 . [^1] In low Earth orbit the driving signal of WEP violation is 500 times stronger than for a torsion balance on the ground PerBraginskyPLA2018
We begin by discussing the principles of a WEP/UFF experiment design in space and the peculiarities of MICROSCOPE (Sec. II). Concentric test masses of different composition are required, approximated by coaxial concentric hollow cylinders. The symmetry axis of each cylinder is the sensitive axis and the displacements along this direction are measured by the variation of capacitance in the bridge formed by each test mass and the cylindrical electrode belts at its ends. The heritage of, and differences from, the GOCE sensors, is discussed. Because of the experiment geometry, the gold wire providing electrical polarization and grounding has to run parallel to the sensitive/symmetry axis. The mechanical stiffness of the wire parallel to this axis is the dominant noise source. The rest position of the mechanical spring does not coincide with the zero of the capacitance bridges and the difference gives rise to a large “constant” force/voltage bias which might vary in the presence of patch charges on the surface of the test cylinders and their electrodes and consequent low frequency fluctuating potentials.
Establishing the nature and source of the noise limiting the measurement is crucial for any experiment and a prerequisite for future improvements (Sec. III).
Systematic effects that were a matter of concern MicroscopeLimitations turned out to be dominated by temperature variations; after calibration and a posteriori correction of all sessions, residual systematic errors were much smaller than random errors (SUEP) or close to them (SUREF) RodriguesSystematicsCQG2022 ; TouboulCQG2022-9 .
The early reports MS-PRL2017 ; CQG2019 showed plots of the square root of the power spectral density (SD) of the random acceleration difference measured in two sessions, and pointed out, in the frequency region of interest, a $1/\sqrt{\nu}$ trend attributed to random thermal noise from internal damping occurring in the gold wires. The final paper MS-PRL2022 does not mention the limiting noise and no longer shows any SD plots. In the batch of companion papers CQGspecialissueMS2022 only a couple of plots are provided TouboulCQG2022-9 which appear to confirm the $1/\sqrt{\nu}$ trend. In general, wire damping is occasionally mentioned but the nature of the noise is not explored further. In one case LiorzouCQG2022-2 a potential minor contribution from electrostatic patch noise is conjectured, with the same frequency dependence.
Assuming random noise from internal damping, and given the fractional acceleration differences provided for the 19 SUEP sessions, one can calculate the SD at the signal frequency, which is a function of the stiffness $k$ (measured) and quality factor $Q$. We find the calculated values of $Q$ exhibit jumps among the different sessions. One such jump was highlighted early on RodriguesBremen2017 , and remained unexplained. We reason $Q$ may change between sessions, as a result of resetting the initial conditions, but not between consecutive data takes without resets. Another noise source, with the same $1/\sqrt{\nu}$ trend and comparable magnitude, must be present.
@@
In GOCE, fluctuating patch potentials with $1/\sqrt{\nu}$ trend at low frequency were one of the main noise sources. We show a fluctuating patch potential at sub-$\mu\rm V$ level at the frequency of the signal would be sufficient to explain the jumps. This is because the fluctuating voltage multiplies a considerably large voltage needed to keep the test mass motionless at an arbitrary chosen zero of the sensor. As a result, a large force bias arises, hence, by way of the fluctuating patch potentials, a fluctuating zero.
The experiment was affected by anomalous acceleration peaks (“glitches”) occurring simultaneously in the four test masses (Sec. IV). The issue has beset all geodesy missions carrying precision accelerometers, except GOCE. It is commonly attributed to release of stress energy accumulated in the multilayer insulation by cycles of heating and cooling induced by the Sun and Earth. GOCE was exempt thanks to an ad hoc, stiffer than usual, solution adopted for the insulation. We find evidence in favor of this interpretation in the description of MICROSCOPE’s glitches, which appear in the frequency domain as double lines, due to the synodic frequencies w.r.t. Sun and Earth, and their harmonics, being very close to one another.
In MICROSCOPE glitches were very numerous, between 14000 and 40000 in 120 orbits. Data intervals at the spikes were removed and replaced with artificially reconstructed data. This procedure does not completely eliminate the effects of glitches, in particular at frequency of the signal, and it is not mandatory. The WEP ground tests by the Eöt-Wash group Adel2008 provide a successful example of data analysis in the presence of outliers carried out without introducing artificial data, exploiting the fact that the frequency and phase of the target signal are known. We argue that the same procedure is applicable to the MICROSCOPE test in orbit, and suggest that –if applied– it would definitively settle the experiment result.
Conclusions are drawn in Sec. V.
II Main features of the MICROSCOPE accelerometers to test WEP/UFF in space
Aside from the inertial force resulting from air drag and solar radiation pressure, which is compensated by drag free control, test masses inside a spacecraft in low Earth orbit are dominated by the effects of Earth tides, which for each TM are linear with the separation from the center of mass (CoM) of the spacecraft while –within the same sensor unit– are proportional to the offset between the CoM of the two test masses.
For these reasons the TMs must be (nominally) concentric and as close as possible to the CoM of the spacecraft, which in MICROSCOPE is constrained by the $17.1\,\rm cm$ physical separation between the two sensor units Touboul2012 . The choice is of coaxial, concentric, hollow test cylinders within each unit. In MICROSCOPE each test cylinder is part of a capacitive accelerometer, designed and built by ONERA, and the sensitive axis is the symmetry axis of the cylinder ($X$ axis).
In most ONERA accelerometers (e.g. in GOCE) as well as in LISA, the TM is a cube (or a parallelepiped) and linear displacements are detected via capacitance variations in the classical “gap-variation” scheme. The TM is located halfway between two capacitance plates facing its two opposite faces perpendicular to the sensitive axis; as the TM moves away from the central “zero” position the capacitance increases on one side and decreases on the other, the variation yielding the displacement that has generated it.
With a hollow cylinder test mass ONERA has opted for the “area-variation” scheme. The electrodes are two conductive belts surrounding the cylinder at its two ends; as the cylinder moves away from the zero position the capacitance variation yields a measurement of the displacement as in the previous case.
@@ II Main features of the MICROSCOPE accelerometers to test WEP/UFF in space
As shown in ONERA1999 , a key difference between the two schemes is that in gap-variation the electrostatic stiffness is negative and usually dominates over the mechanical, positive stiffness of the gold wire, hence the TM is unstable. Instead, with area-variation the electrostatic stiffness is smaller and positive, like the stiffness of the wire, and therefore the TM is stable. In gap-variation the capacitance is inversely proportional to the gap squared, while in area-variation it is inversely proportional to the gap, and smaller gaps are known to give rise to higher disturbances. Instead of reducing the gap a larger surface of the electrodes is used.
@@
In the MICROSCOPE accelerometers the gold wire is parallel to the symmetry/sensitive $X$ axis. Thus, as the TM moves along this axis the wire is subjected to a longitudinal force and responds to it with a stiffness $k_{\parallel}$; instead, in the perpendicular $Y,\,Z$ directions the force is perpendicular to the wire which responds with a different stiffness $k_{\perp}$.
@@
The difference between $k_{\parallel}$ and $k_{\perp}$ is well illustrated in GOCE. In that case the TM of each accelerometer is a parallelepiped with two “ultra sensitive” axes ($Y,\,Z$) while the third one ($X$) is less sensitive. It is based on gap-variation and the gold wire ($5\,\rm\mu m$ diameter and $2.5\,\rm cm$ length) is directed along the $X$ axis. The stiffnesses reported are $k_{\parallel}=10^{-3}\,\rm N/m$ along $X$ and $k_{\perp}=1.8\times 10^{-6}\,\rm N/m$ along $Y,\,Z$; the gaps are of $32\,\rm\mu m$ along $X$ and $300\,\rm\mu m$ along $Y,\,Z$, and the larger face of the parallelepiped ($4\times 1\,\rm cm^{2}$) is perpendicular to the $X$ axis while the $Y,\,Z$ axes are perpendicular to the smaller faces (of $1\times 1\,\rm cm^{2}$) E2ESimulatorGOCE . The $X$ axis was used to test the accelerometer in the lab, where the TM must be levitated against the local gravitational acceleration instead of the much smaller tidal acceleration in orbit.
@@
The gold wire of MICROSCOPE’s test cylinders has the same length and only a slightly bigger diameter ($7\,\rm\mu m$ instead of $5$) than the wire of GOCE, hence the stiffness along the $X$ axis must be very close to $k_{\parallel}=10^{-3}\,\rm N/m$ reported for GOCE, and indeed the in orbit measurements found values around $\simeq 10^{-3}\,\rm N/m$ for all four test cylindersCQG2019 ; ChhunCQG2022-5 .
@@
Instead BergeCQG2022-10 , the stiffness expected for the MICROSCOPE wire was based on the theoretical formula:
@@
The reason for the large discrepancy is that the theoretical stiffness (1), as well as the experimental tests supporting it WT-RSI2000b , apply to the response of the wire to a force acting perpendicularly to it, i.e. $k_{w}=k_{\perp}$, whereas the relevant stiffness is that acting along the symmetry/sensitive axis, $k_{\parallel}$.
@@
Within the mission team, the in orbit measurements ChhunCQG2022-5 have been questioned by BergeCQG2022-10 , Sec. 5.1. However, the model used is incorrect, giving –for a wire along the $X$ axis– zero stiffness in the $Y,Z$ directions instead of the correct value resulting from (1).

### v4

# On the first test of the Weak Equivalence Principle in low Earth orbit

@@ Abstract
The Weak Equivalence Principle is the founding pillar of General Relativity and as such should be verified as precisely as possible. The Microscope experiment tested it in low-Earth orbit, finding that Pt and Ti test masses fall toward Earth with the same acceleration to about 1 part in $10^{15}$, an improvement of about two orders of magnitude over ground tests. Space missions, even if small, are expensive and hard to replicate; yet, the essence of physics is repeatability. This work is an assessment of the Microscope results based on the laws of physics and knowledge from previous experiments, focusing on the limiting thermal noise and the treatment of acceleration outliers. Thermal noise reveals anomalies that we explain by stray sub-$\mu\rm V$ potentials caused by patch charges, giving rise to an unstable zero. The measurements were affected by numerous acceleration spikes occurring at the synodic frequencies relative to the Earth (the signal frequency) and the Sun, which we interpret as evidence of a thermal origin. In Microscope authors’ analysis, the spikes were removed and the resulting gaps replaced with artificial data (up to $35\%$, $40\%$ of the sessions data), which retain memory of the gaps and may simulate or cancel an effect (signal or systematic). An alternative approach based exclusively on real measured data would avoid any ambiguity. The lessons of Microscope are crucial to any future improved mission.
@@ I Introduction
The Weak Equivalence Principle (WEP), or rather not a “principle” but the “hypothesis” of a complete physical equivalence of the gravitational field and the corresponding acceleration of the reference frame –as Einstein expressed it in 1907– is the founding pillar of General Relativity. In Newton’s formulation it was the equivalence of inertial to gravitational mass and the Universality of Free Fall (UFF).1 As such, it should be tested as precisely and accurately as possible, something physicists have been doing since Galileo’s time.
@@
Even a small space mission like Microscope ($300\,\rm kg$) is a “big science” project hard to replicate. Yet, the essence of science is repeatability. The mission was carried out by a single team as a single experiment. The measurement data are publicly available, but the practice of having two independent teams performing the data analysis in parallel (e.g. as in the Hipparcos mission of ESA) has not been pursued. On the other hand, a new data analysis performed post facto, without the in depth knowledge (of the experimental setup, spacecraft, software and operations) that comes from first hand participation in a satellite mission, would be highly controversial and ultimately inconclusive –if it ended up with anything inconsistent with the original results– or irrelevant –if it followed too closely the steps of the first analysis.
Does this mean that the result of Microscope shall be fully assessed only by another space mission of comparable or higher sensitivity? The answer is no, because it is possible to break down the experiment into its key features, each of which can be investigated on the basis of basic physics arguments and well established knowledge, both theoretical and experimental.
The basic features of the Microscope experiment are addressed in Secs. II and III. Sec. II is concerned with the test masses, the accelerometers and the acceleration measurements, a potential violation signal from Earth, its signature and how it would be measured by a Sensor Unit (SU). In particular we show the $500$ gain factor in space and stress the difference between an inherently differential instrument like the torsion balance (no violation, no signal) and the SU of Microscope in which the acceleration difference is reconstructed from the individual measurements of two independent accelerometers. The relevance and consequences of such feature are addressed in Secs. III and IV.
Sec. III deals with the forces acting on the test masses, starting from drag and drag-free control and focusing on the electrostatic control of the test mass in the presence of Earth tidal forces, mechanical stiffness, biases, patch charges and fluctuating patch potentials.
@@
Establishing the nature and source of the noise limiting the measurement is crucial for any experiment and a prerequisite for future improvements. In Sec. IV we analyze the limiting random noise. Since the preliminary results 5 the residual acceleration difference was reported to be limited by thermal noise with $1/\sqrt{\nu}$ frequency dependence attributed to internal damping in the gold wires. However, our analysis of the final results for all measurement sessions 4; 6 reveals anomalies which are inexplicable within the current best knowledge of this noise. In GOCE fluctuating patch potentials with $1/\sqrt{\nu}$ trend at low frequency were one of the main noise sources. We show that such a potential at sub-$\mu\rm V$ level at the frequency of the signal would be sufficient to explain the observed anomalies. This is because the fluctuating voltage multiplies a considerably large voltage needed to keep the test mass motionless at an arbitrary chosen zero of the sensor. As a result, a large force bias arises, hence, by way of the fluctuating patch potentials, a fluctuating zero. An unstable zero is a known issue in WEP tests, related to the problem of initial conditions (or release) errors, affecting all tests not designed as intrinsically null experiments.
The presence of patch charges and consequent fluctuating potentials at the frequency of the signal should be revisited in order to firmly establish its contribution to the limiting noise, starting from the knowledge acquired with GOCE.
Systematic effects that were a matter of concern,7 turned out to be dominated by temperature variations. After calibration of the thermal response and a posteriori correction, residual systematic errors were much smaller than random errors or close to them.9; 8 This was achieved thanks to an effective multilayer thermal insulation blanket (MLI). However, MLI blankets are the likely source of large acceleration spikes (“glitches”) occurring simultaneously in the four test masses (Sec. V). GOCE was exempt thanks to an ad hoc, stiffer than usual, solution adopted for the insulation.
Microscope’s glitches appear to be induced by thermal stress from the Earth and the Sun on the MLI. They occur at the synodic frequency relative to Earth (also the frequency of a violation signal) and the Sun, their harmonics and their difference. Outliers were removed and the ensuing gaps (up to $35\%$, $40\%$ of the sessions data) were filled with artificial reconstructed data. However, gaps and artificial data retain memory of the outliers they have replaced and may therefore mimic a violation signal or cancel an effect (signal or systematic).
An alternative approach relying solely on real measured data has been successfully used in the Eöt-Wash ground tests, with only $7\%$ of data removed.2 We argue that the same procedure is applicable to the Microscope test in orbit.
Conclusions are drawn in Sec. VI, along with the main takeaway lessons for future improvements.
II The accelerometers of Microscope and the signature of a potential WEP violation signal
Inside an Earth orbiting spacecraft its center of mass (CoM) is the only zero force point (perfect weightlessness) since in this point the Earth’s gravitational attraction is perfectly balanced by the inertial centrifugal force arising in the non inertial reference frame of the spacecraft. Any Test Mass (TM) located at a non zero position vector relative to the CoM of the spacecraft is subjected to a tidal acceleration because the gravitational field is not uniform, resulting in gravity gradient forces referred to as tidal forces (since ancient times sea tides were attributed to the Moon and the Sun because of their typical frequencies).
If a TM inside an Earth orbiting spacecraft violates the equivalence between its inertial and gravitational mass ($m_{1}^{i}$ and $m_{1}^{g}$ respectively) it is $m_{1}^{g}=m_{1}^{i}(1+\eta_{{}_{1}})$ where $\eta_{1}\neq 0$ is a dimensionless parameter quantifying the violation. In this case the Earth acts on the TM not only through the tidal acceleration, but also with an additional acceleration $a_{{}_{1}}=g\eta_{{}_{1}}$ where $g=GM_{\oplus}/r^{2}$ is the monopole gravitational attraction from the Earth (with $M_{\oplus}$ the mass of the Earth, $r$ the orbital distance of the spacecraft, $G$ the universal constant of gravity). If another TM of different composition violates the equivalence of inertial to gravitational mass at level $\eta_{{}_{2}}$, and the two masses are perfectly concentric (hence they are affected by the same tidal acceleration), there will be a differential acceleration between the two given by:
$a=a_{{}_{2}}-a_{{}_{1}}=g\eta\ \ \ \ \ ,\ \ \ \ \eta=\eta_{{}_{2}}-\eta_{{}_{1}}\ \ .$ (1)
[…]
The concepts of equivalence between inertial and gravitational mass, UFF and WEP can all be used to mean the same thing.1 The theoretical framework is that of a 2-body problem (TM in the field of the source body) with WEP violation, and the violation signal lies in the orbit plane of the spacecraft.10 Since the sensor is rigid with the spacecraft it is natural to work in the non inertial reference frame of the spacecraft centered on its CoM relative to which –in case of space fixed attitude– the Earth orbits at angular velocity of modulus $\omega_{orb}=GM_{\oplus}/r^{3}$. Since the signal has two degrees of freedom it is desirable that the sensor too be sensitive in 2D (the orbit plane). The advantages over 1D sensors is to double the measurement data (and thus double the integration time $t_{int}$ and reduce the random noise by $\sqrt{2}$, since it is proportional to $1/\sqrt{t_{int}}$), and to avoid the risk of noise leakage from the other degree of freedom which is less sensitive.
From Eq.(1) it is apparent that, for a given sensitivity of the experiment to differential accelerations between the TMs, the larger the driving signal, the smaller the resulting value of $\eta$, the better the WEP test.
At the orbiting altitude of Microscope ($\simeq 700\,\rm km$) it is $g=7.9\,\rm ms^{-2}$. For test masses suspended on the ground, as in a pendulum or in torsion balances, the driving signal is given by the inertial centrifugal acceleration $\omega_{\oplus}^{2}R_{\oplus}\cos{\vartheta}\sin{\vartheta}$ due to the diurnal rotation of the Earth $\omega_{\oplus}=2\pi/86164\,\rm rad/s$ which balances the component of the gravitational attraction of the Earth in the horizontal plane tangent to its surface ($R_{\oplus}\simeq 6.38\times 10^{6}\,\rm m$ the Earth radius) at a given latitude $\vartheta$. Its maximum value (at $45^{\circ}$ latitude) is $\simeq 0.0169\,\rm ms^{-2}$.
[…]
Any non zero offset between the centers of mass of the TMs results in a differential tidal acceleration directly proportional to the offset. In order to reduce the offset, and also for practical reasons, the TMs of a WEP test in space are designed as two concentric, coaxial, hollow cylinders.
In Microscope each test cylinder forms an independent accelerometer; it is suspended electrostatically, its symmetry axis $X$ is the best sensitive axis, it lies in the $X,Z$ orbit plane and it is devoted to the detection of a potential violation signal. As far as WEP testing is concerned it is therefore a 1D sensor. However, each test cylinder has 6 degrees of freedom, and must be controlled in all of them. In this work we focus on the $X$ sensitive axis.
Two coaxial accelerometers form what is called a Sensor Unit (SU), and the spacecraft carries two of them (Fig. 1), one with different composition $\rm Pt$-$\rm Ti$ test cylinders for testing the WEP, named SUEP, and the other with $\rm Pt$-$\rm Pt$ cylinders named SUREF as a reference sensor because equal composition masses cannot violate the WEP even if their masses are different.
In each accelerometer two conductive electrode belts are placed at the two ends of the cylinder and measure the capacitance variations along the symmetry $X$ axis according to the “area-variation” scheme (Fig. 2).
A tiny gold wire parallel to the symmetry axis connects the test cylinder to the cage enclosing it (rigid with the spacecraft) for electric grounding and polarization of the test mass.
Like in all accelerometers by ONERA, the test masses of Microscope –although in the shape of hollow cylinders rather than cubes or parallelepipeds– are not allowed to move. A control voltage is applied to nullify any displacement so as to keep the TM motionless (except for the residual noise that cannot be controlled) at a chosen capacitive zero $O_{C}$ (which depends on geometry, construction, mounting and electronics). By calibration, a scale factor is established (in $\rm ms^{-2}/V$) by which the applied electric potential yields the acceleration acting on the test cylinder. Then, in Microscope a scale matching is performed between the two accelerometers of each SU and their acceleration difference is obtained, which is the physical quantity of interest for testing the WEP. This shows that the SU is not an intrinsically differential sensor, meaning that it does not measure the differential acceleration of the test masses directly, but provides it as the difference of two separate acceleration measurements (Secs. III and IV).
Figure 1: Microscope Sensor Units in space (left, figure taken from Microscope presentations) and WEP violation signal at zero spin (right). The orbit ($\simeq 700\,\rm km$ altitude) is near-circular ($e\simeq 10^{-4}$) and near-polar ($i\simeq 98.2^{\circ}$, making it sun-synchronous with solar panels facing the Sun, and the descending node at dawn). In the reference frame of the spacecraft, the $Z$ axis points at the CoM of the Earth and the $Y$ axis points in the direction of the outward orbit normal. The satellite spins clockwise (opposite to the orbital motion) so that the signal frequency is the sum of the orbit and spin rate. The separation between the centers of mass of the two Sensor Units is along the $Y$ axis, so that, projected on the $X,Z$ orbit plane, the position of the centers of mass of the test cylinders can be closer to the CoM of the satellite, as constrained by construction and mounting errors. At zero spin a WEP violation signal would act along the symmetry axis at $\omega_{orb}$ (right sketch, section of SUEP sensor on the orbit plane with the different composition cylinders depicted in different colors).
Figure 2: Microscope test cylinder with $X$ axis electrodes (Fig.5, Ref.(12)). In each accelerometer two conductive electrode belts are placed at the two ends of the cylinder with a gap of $598\,\mu\rm m$. If the CoM of the cylinder moves away from the central position, the area of the cylinder facing its electrode increases on one side and decreases on the other (“area-variation”), and so does the capacitance, by an amount depending on the displacement, hence on the acceleration that generated it.
If the spacecraft attitude is fixed relative to inertial space the signal of a WEP violation from the Earth on the sensitive axis of SUEP sensor is an acceleration difference at the orbital frequency and phase as shown in Fig. 1, right sketch, of unknown amplitude and sign. If the spacecraft spins relative to inertial space around the $Y$ axis perpendicular to the orbit plane as shown in Fig. 1, a potential violation signal is up-converted to the higher orbital plus spin frequency (the signs are opposite). Of course, any other effect at the orbital frequency, such as drag, is up-converted too.
The advantage of rotation is to reduce the noise which is known to decrease at higher frequency (Sec. IV) and also to average thermal effects. However, as shown in both sketches of Fig. 1, by the very design of the Microscope sensor, rotation cannot occur around the symmetry axis which would ensure a higher spin rate and passive rotation by conservation of angular momentum, thus limiting attitude maneuvers only to keep the solar panels close to the perpendicular to the Sun.
III Forces acting on Microscope accelerometers and their control
Residual air along the orbit and solar radiation impinging on the outer surface of the spacecraft give rise to a force on its CoM we will globally refer to as “drag”. As a result, the reference frame of the spacecraft in which measurements are performed is an accelerated non inertial frame and any TM inside it is subjected to an inertial acceleration equal and opposite to the non gravitational acceleration of the CoM of the spacecraft.
Ideally, being the same on all TMs (common mode effect) it should not affect the differential acceleration between the two TMs of the sensor unit. In reality, the two accelerometers of the SU do not respond with exactly the same acceleration. Only a limited rejection of common mode effects is achieved by scale matching in orbit. The residual differential effect of drag has the same frequency as the target signal and almost $90^{\circ}$ phase difference (at Microscope’s low altitude residual air dominates over solar radiation and the effect of drag is mostly along track, while a violation signal would point towards or away from Earth). A remaining differential component –after drag free control and partial rejection– with the same frequency and phase as the signal would mimic a violation signal; it is therefore necessary to ensure that the entire residual differential effect be smaller than the target.
The drag free control system receives the non-gravitational acceleration of the spacecraft measured by the accelerometers and synthesizes the commands to the actuators (thrusters) that result in equal forces of opposite sign (with limited precision, in a limited control bandwidth). In general, one of the two SUs is used to drive the drag free control while the other takes science data.
In Microscope the acceleration due to drag is about $10^{-8}\,\rm ms^{-2}$. The requirement placed on drag free control is to partially compensate it leaving a residual common mode acceleration of about $10^{-12}\,\rm ms^{-2}$. This would be then partially rejected by matching in orbit the scale factors of the two accelerometers of each SU such that the residual acceleration difference caused by drag is reduced to $3$-$4\times 10^{-15}\,\rm ms^{-2}$, with a WEP violation target of $7.9\times 10^{-15}\,\rm ms^{-2}$.
The test cylinders of Microscope’s accelerometers are not free masses. They are controlled electrostatically, and they are subject to electrostatic stiffness and to the mechanical stiffness of the gold wire.
@@ III Forces acting on Microscope accelerometers and their control
In most ONERA accelerometers (e.g. in GOCE) as well as in LISA, the TM is a cube (or a parallelepiped) and displacements are detected via capacitance variations in the classical “gap-variation” scheme in which the TM is located halfway between two capacitance plates facing its two opposite faces perpendicular to the sensitive axis. This scheme is used in Microscope for the $Y,Z$ axes, while the $X$ symmetry/sensitive axis is based on area-variation. As shown in Ref.(13), in gap-variation the electrostatic stiffness is negative and usually dominates over the mechanical, positive stiffness of the gold wire, hence the TM is unstable. Instead, in the area-variation scheme the electrostatic stiffness is smaller and positive, like the stiffness of the wire, and the TM is stable. In gap-variation the capacitance is inversely proportional to the gap squared, while in area-variation it is inversely proportional to the gap, but smaller gaps are known to give rise to higher disturbances. Instead of reducing the gap a larger surface of the electrodes is used to increase the capacity.
@@
In Microscope, as a test cylinder moves along the $X$ axis the gold wire is subjected to a longitudinal force and responds to it with a stiffness $k_{\parallel}$; instead, in the perpendicular $Y,Z$ directions the force is perpendicular to the wire which responds with a different stiffness $k_{\perp}$.
@@
The difference between $k_{\parallel}$ and $k_{\perp}$ is well illustrated in GOCE. In that case the TM of each accelerometer is a parallelepiped with two “ultra sensitive” axes ($Y,Z$) while the third one ($X$) is less sensitive. It is based on gap-variation and the gold wire ($5\,\rm\mu m$ diameter and $2.5\,\rm cm$ length) is directed along the $X$ axis. The stiffnesses reported are $k_{\parallel}=10^{-3}\,\rm N/m$ along $X$ and $k_{\perp}=1.8\times 10^{-6}\,\rm N/m$ along $Y,\,Z$; the gaps are of $32\,\rm\mu m$ along $X$ and $300\,\rm\mu m$ along $Y,\,Z$, and the larger face of the parallelepiped ($4\times 1\,\rm cm^{2}$) is perpendicular to the $X$ axis while the $Y,\,Z$ axes are perpendicular to the smaller faces (of $1\times 1\,\rm cm^{2}$).14 The $X$ axis was used to test the accelerometer in the lab, where the TM must be levitated against the local gravitational acceleration instead of the much smaller tidal acceleration in orbit.
@@
The gold wire of Microscope’s cylinders has the same length and only a slightly bigger diameter than the wire of GOCE ($7\,\rm\mu m$ instead of $5$), hence the stiffness along the $X$ axis must be very close to $k_{\parallel}=10^{-3}\,\rm N/m$ reported for GOCE, and indeed the in orbit measurements found values around $\simeq 10^{-3}\,\rm N/m$ for all four test cylinders.12; 15
@@
Instead,16 the stiffness expected for the Microscope wire was based on the theoretical formula:
@@
The reason for such a large discrepancy is that the theoretical stiffness of Eq.(2), as well as the experimental tests supporting it,17 apply to the response of the wire to a force acting perpendicularly to it, i.e. $k_{w}=k_{\perp}$, whereas the relevant stiffness is that acting along the symmetry/sensitive axis, $k_{\parallel}$.
@@
Within the mission team, the in orbit measurements 15 have been questioned by Ref.(16), Sec. 5.1. However, the model used is incorrect, giving –for a wire along the $X$ axis– zero stiffness in the $Y,Z$ directions instead of the correct value resulting from Eq.(2).